3x divided by 4 equals -8
Answer by checkley71(8403)

3x/4=-8
3x=-32
x=-32/3
x=-10 2/3 solution
proof
3(-10 2/3)/4=-8
3(-32/3)/4=-8
(-96/3)/4=-8
(-96/3*1/4)=-8
-96/12=-8
-8=-8